OpenStudy (anonymous):

Evaluate C(7, 7). a)0 b)1 b)7

OpenStudy (anonymous):

The formula here is \[C(n,m)=\frac{n!}{m!(n-m)!}.\]

OpenStudy (anonymous):

i got 0 ?

OpenStudy (anonymous):

\(0!\) is equal to \(1.\) (It's weird, I know.)

OpenStudy (anonymous):

Wait so 0 or 1 ???

OpenStudy (anonymous):

What do you think? (not being rude, I'm curious which you think.)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Correct. An interesting fact is \(C(k,k)=\frac{k!}{k!(k-k)!}=\frac{k!}{k!0!}=\frac{k!}{k!}=1\) for any \(k.\)